OpenStudy (anonymous):

A rectangle has an area of 72 square units. The width of the rectangle is 9 units. The length of the rectangle is 2x + 4. What is the rectangle's length? 6 7 8 9

OpenStudy (phi):

what is the formula for the area of a rectangle ?

OpenStudy (anonymous):

lenth over width?

OpenStudy (phi):

Length * width = area you know width, and you know area. if you put in the numbers what do you get ?

OpenStudy (anonymous):

8

OpenStudy (phi):

L * W = A replace W with 9 and A with 72 to get L*9 = 72 divide both sides by 9 L*9/9 = 72/9 L = 8